Aqueous solution of uria of freezing point is 0.4°c. What will be boiling point this solution (Kf=3.632Kb)

Grade:12

1 Answers

Vikas TU
7 years ago
Tf = iKfm
0.4 = 1*3.632Kb*m
 
Boiling Point,
Tb = 1*Kb*m
Dividing both eqns. we get,
0.4/Tb = 3.632
Tb = 0.4/0.3632 => 1.10132 ‘C